The sum of two numbers is twenty-four. Four less than three times the smaller is twelve less than twice the larger. Find the two numbers.
Answer by lwsshak3(11628)

The sum of two numbers is twenty-four. Four less than three times the smaller is twelve less than twice the larger. Find the two numbers.
..
let x=larger number
then 24-x=smaller number
3(24-x)-4=2x-12
72-3x-4=2x-12
5x=72-4+12=80
x=16
ans:
larger number=16
smaller number=24-16=8